# Probing the cosmic sterile-neutrino background with IceCube

### v1

@@ Abstract
In this paper, we take a close look at the interaction between the TeV–PeV energy astrophysical neutrinos and a hypothetical cosmic sterile-neutrino background. These interactions yield absorption features, also called “dips”, in the astrophysical neutrino spectrum, which are studied using the deposited energy distribution of high-energy starting events (HESE) in the IceCube detector. We improve upon the previous analysis by including the effects of regeneration and a realistic source distribution on the propagation of astrophysical neutrinos. We use the latest 7.5-year HESE dataset and include the observation of Glashow resonance in our analysis. We evaluate the impact of these dips on the inferred spectral index and overall normalization of the astrophysical neutrinos. We find a mild preference for dips in the 300–800 TeV range, and the best-fit parameters for the mass of sterile-neutrino and the mediator are 0.5 eV and 23 MeV, respectively. We find that the inclusion of these absorption features lowers the spectral index of astrophysical neutrinos to $2.60^{+0.19}_{-0.16}$. The lower spectral index can reduce the disagreement with the Northern Tracks sample but requires dedicated analysis. We also forecast the event spectrum for IceCube-Gen2 for the two different fits.
@@ I Introduction
The IceCube neutrino observatory has detected neutrinos of astrophysical origin in the TeV–PeV range in several detection channels. The latest 7.5 years (2635 days) high energy starting events (HESE) sample consists of 102 events, of which 60 events have reconstructed deposited energy above 60 TeV Abbasi et al. (2021). The IceCube experiment has also observed a Glashow Resonance (GR) event Glashow (1960) in the 4.6 years PeV energy partially contained events (PEPE) sample Aartsen et al. (2021a). IceCube also maintains a catalog of through-going muon-neutrinos from the northern hemisphere (henceforth called Northern Tracks). A large fraction of the Northern Tracks are below 100 TeV and originate via atmospheric neutrinos. The highest energy events from this sample have astrophysical origin.
It is abundantly clear that these neutrinos have an extra-galactic origin and are likely to originate in sources capable of accelerating cosmic rays to EeV-scale energies. The neutrinos are produced in subsequent interactions of these ultra-high energy cosmic rays with ambient gas or radiation. At the time of writing, only a handful of such sources have been confirmed Aartsen et al. (2018a, b, 2020); Abbasi et al. (2022a). If the neutrinos are produced via interactions with ambient gas, i.e. in proton-proton ($pp$) interactions, the neutrinos are expected to have a power-law spectrum Margolis et al. (1978); Stecker (1979); Kelner et al. (2006). In the baseline scenario, the neutrinos oscillate and reach flavor equilibrium before they arrive to the Earth Lipari et al. (2007); Choubey and Rodejohann (2009); Mena et al. (2014, 2014); Aartsen et al. (2015). One usually assumes that an isotropic flux of astrophysical neutrinos is incident on the Earth with a total differential all-flavor neutrino-plus-anti-neutrino spectrum given by
@@
Are the events observed by IceCube consistent with the assumption of SPL? To a large extent, yes. However, there are some tantalizing discrepancies that have motivated a large body of work in the last decade. Some of the explanations are based on dark matter Feldstein et al. (2013); Esmaili and Serpico (2013); Bai et al. (2016); Fong et al. (2015); Rott et al. (2015); Esmaili et al. (2014); Murase et al. (2015); Anchordoqui et al. (2015); Bhattacharya et al. (2017a, b); Hiroshima et al. (2018); Chakravarty et al. (2020), resonances beyond standard model Anchordoqui et al. (2006); Barger and Keung (2013); Dutta et al. (2015); Dey and Mohanty (2016); Dev et al. (2016); Mileo et al. (2016); Chauhan et al. (2018); Dey et al. (2018); Bečirević et al. (2018); Babu et al. (2020); Dey et al. (2021), or non-standard/secret neutrino interactions Ng and Beacom (2014); Ioka and Murase (2014); Ibe and Kaneta (2014); Blum et al. (2014); Araki et al. (2015); Cherry et al. (2014); Kamada and Yu (2015); DiFranzo and Hooper (2015); Araki et al. (2016); Rasmussen et al. (2017); Jeong et al. (2018); Mohanty et al. (2019); Pandey et al. (2019a, b); Bustamante et al. (2020); Creque-Sarbinowski et al. (2021); Mazumdar et al. (2022); Fiorillo et al. (2020a, b); Carpio et al. (2023); Esteban et al. (2021). One such discrepancy is the inferred value of the spectral index in two independent datasets,
@@
We begin with the ansatz that including “dips” in the neutrino flux will result in a lower spectral index from the HESE sample and thus relax the tension. Such dips are a consequence of interaction between astrophysical neutrinos and cosmic background. If the interaction is through a s-channel process, then the absorption feature is narrowly located around the resonance energy $E_{\nu}^{\rm res}=M_{X}^{2}/2m_{b}$, where $M_{X}$ is the mass of the mediator and $m_{b}$ is the mass of the background particle. If the interaction is t-channel, then a broad spectrum feature is expected Mohanty et al. (2019). In Ref. Esteban et al. (2021), the effect of non-resonant contribution to the process was found to be significant for larger couplings, and we have included them in our analysis. With only standard model interactions, the dip will occur for neutrinos with energy EeV Weiler (1982); Brdar et al. (2022). However, if there are non-standard/secret interactions mediated by a MeV-scale boson, then the dips can occur in the TeV–PeV range Ng and Beacom (2014). In this paper, we study the scenario proposed in Ref. Chauhan and Mohanty (2018) where the dip is caused by interactions between astrophysical neutrinos and a cosmic sterile-neutrino background.
Figure 1: The test statistic (see Equation 5) for the single power-law fit of 7.5 years HESE sample Abbasi et al. (2021) is shown in blue. In orange, we show the test statistic after including absorption features and Glashow resonance. The horizontal grey dashed lines are used to denote the 68% C.L. and 90% C.L. intervals. We have shown the results from the 11-year Northern Tracks sample Stettner (2020) in green.
The light sterile neutrino is a well-motivated extension of the standard model and has been the subject of investigation in a number of experiments. Currently, there is evidence in support of  Anselmann et al. (1995); Hampel et al. (1998); Aguilar-Arevalo et al. (2009, 2010, 2001); Mention et al. (2011); Aguilar-Arevalo et al. (2018); Serebrov et al. (2021) as well as against Abratenko et al. (2023) the sterile neutrino. For a recent review, we refer to Ref. Dasgupta and Kopp (2021). The presence of a fully thermalized sterile neutrino is in conflict with the measurement of $N_{\text{eff}}$ from cosmological probes Hamann et al. (2011); Cooke et al. (2018); Aghanim et al. (2020). If the sterile neutrino has self-interactions via a MeV scale vector mediator, their recoupling can be delayed, and the $N_{\text{eff}}$ constraints can be avoided Hannestad et al. (2014); Dasgupta and Kopp (2014). The case for pseudo-scalar mediator is studied in Ref. Archidiacono et al. (2015). Despite the early success, the self-interacting sterile neutrino scenario has been tested against many cosmological observations Saviano et al. (2014); Mirizzi et al. (2015); Forastieri et al. (2017); Choi et al. (2018); Song et al. (2018). It was shown in Ref. Chu et al. (2018) that the basic scenario is in conflict with bounds on the sum of neutrino mass, and some resolutions have been provided as well. One must note that these results depend on the choice of background cosmology as $\Lambda$CDM, and the bounds will change if other scenarios are considered Roy Choudhury and Naskar (2019); Roy Choudhury and Choubey (2018); Roy Choudhury and Hannestad (2020). Recent analyses of large-scale structure data have also indicated the preference for self-interacting neutrinos He et al. (2024); Camarena et al. (2023). With this as motivation, we ask if the dip due to a cosmic sterile-neutrino background can be probed in IceCube and whether the inclusion of these features leads to an appreciable decrease in the inferred spectral index. The main result of this paper is shown in Figure 1. We find that the systematic inclusion of the dips results in,
[…]
which is in better agreement with the Northern Tracks sample. As discussed below, the errors are underestimated due to the limitations of our analysis. For a quantitative assessment of the relaxation in the tension between the two datasets, one must evaluate the impact of dips in the spectrum on the Northern Tracks sample. However, this exercise is beyond the scope of our paper. We have shown that an ‘SPL with dips’ is a viable fit to the present HESE sample and results in a lower spectral index than an SPL-only explanation. One must note that our conclusion also extends to other origins of the dip(s).
The paper is organized as follows. In Section II, we discuss the model, propagation of astrophysical neutrinos, our estimate of the HESE event rate, and the likelihood. In Section III, we provide the main results and end with a short summary in Section IV.
@@ II Model description and methods
In this work, we follow the model outlined in Ref. Dasgupta and Kopp (2014), which is briefly described below. The sterile neutrinos are assumed to have self-interactions mediated by a new MeV-scale vector boson. The interaction Lagrangian is given by,
$-\mathcal{L}_{\rm int}=g_{X}\bar{\nu}_{s}\gamma_{\mu}\frac{1}{2}(1-\gamma_{5})\nu_{s}X^{\mu}\,,$ (2)
@@
where $X^{\mu}$ is the mediator and $g_{X}$ is the self-interaction coupling constant. In this work, we fix the coupling strength as $g_{X}=0.1$, which is sufficiently large for the re-coupling of sterile neutrinos Chu et al. (2015). The mass-eigenstates and the flavor-eigenstates are related by a $4\times 4$ rotation matrix characterized by six mixing angles. The three mixing angles from the active sector are fixed to their measured values from Ref. Esteban et al. (2020), and the remaining three angles are free parameters of the model. Due to this mixing, all four mass-eigenstates interact via the new mediator, and interactions between propagating mass-eigenstates and background neutrinos are possible Chauhan and Mohanty (2018). In this simplified model, we have five new physics parameters – three mixing angles ($\theta_{14}\,,\theta_{24}\,,\theta_{34}$), mass of the sterile neutrino ($m_{s}$), and mass of the mediator ($M_{X}$). In a UV-complete model, additional degrees of freedom may be required for generating the mixing and mass of the gauge boson and for anomaly cancellation. The alternative model with a pseudoscalar mediator is cosmologically more viable Dasgupta and Kopp (2021), but has additional constraints from meson decays Fiorillo et al. (2020a). We anticipate that similar bounds will exist on vector mediators, but their evaluation is beyond our scope. Our main conclusion, i.e., the spectral index is lowered because of the dips, does not depend on the details of the model causing the dips in the spectrum.
@@
[…]
@@
In this equation, the first term on the right-hand side accounts for the cosmological expansion. The second term, called the source term, accounts for the production of high-energy neutrinos due to astrophysical sources. In this work, we assume that the TeV–PeV neutrinos originate in $pp$-sources and following Ref. Esteban et al. (2021), the source term is taken to be a single power law with normalization proportional to the star-formation rate. The third term accounts for the interaction with background neutrinos, and the last term accounts for the regeneration of lower-energy neutrinos after interactions. These equations have been solved using the publicly available code nuSIprop Esteban et al. (2021), which solves the Boltzmann equations for standard three-flavor neutrinos in the presence of non-standard interaction. We have extended the code to include the sterile neutrino. These are improvements on the work done in Ref. Chauhan and Mohanty (2018) where the source term was ad-hoc, and regeneration was ignored.

### v2

@@ Abstract
In this paper, we take a close look at the interaction between the TeV–PeV energy astrophysical neutrinos and a hypothetical cosmic sterile-neutrino background. These interactions yield absorption features, also called “dips”, in the astrophysical neutrino spectrum, which are studied using the deposited energy distribution of high-energy starting events (HESE) in the IceCube detector. We improve upon the previous analysis by including the effects of regeneration and a realistic source distribution on the propagation of astrophysical neutrinos. We use the latest 7.5-year HESE dataset and include the observation of Glashow resonance in our analysis. We evaluate the impact of these dips on the inferred spectral index and overall normalization of the astrophysical neutrinos. We find a mild preference for dips in the 300–800 TeV range, and the best-fit parameters for the mass of sterile-neutrino and the mediator are 0.5 eV and 23 MeV, respectively. We find that the inclusion of these absorption features lowers the spectral index of astrophysical neutrinos to $2.60^{+0.19}_{-0.16}$. We show qualitatively that the lower spectral index from HESE sample can reduce the disagreement with the Northern Tracks sample. We also forecast the event spectrum for IceCube-Gen2 for the two different fits.
@@ I Introduction
The IceCube neutrino observatory has detected neutrinos of astrophysical origin in the TeV–PeV range in several detection channels. The latest 7.5 years (2635 days) high energy starting events (HESE) sample consists of 102 events, of which 60 events have reconstructed deposited energy above 60 TeV Abbasi et al. (2021a). The IceCube experiment has also observed a Glashow Resonance (GR) event Glashow (1960) in the 4.6 years PeV energy partially contained events (PEPE) sample Aartsen et al. (2021a). IceCube also maintains a catalog of muon-like tracks originating from the northern hemisphere (henceforth called Northern Tracks). A large fraction of the Northern Tracks are below 100 TeV and originate via atmospheric neutrinos. The highest energy events from this sample have astrophysical origin.
It is abundantly clear that these high-energy neutrinos have an extra-galactic origin and are likely to originate in sources capable of accelerating cosmic rays to EeV-scale energies. The neutrinos are produced in subsequent interactions of these ultra-high energy cosmic rays with ambient gas or radiation. At the time of writing, only a handful of such sources have been confirmed Aartsen et al. (2018a, b, 2020); Abbasi et al. (2022a). If the neutrinos are produced via interactions with ambient gas, i.e. in proton-proton ($pp$) interactions, the neutrinos are expected to have a power-law spectrum Margolis et al. (1978); Stecker (1979); Kelner et al. (2006). In the baseline scenario, the neutrinos oscillate and reach flavor equilibrium before they arrive to the Earth Lipari et al. (2007); Choubey and Rodejohann (2009); Mena et al. (2014, 2014); Aartsen et al. (2015). One usually assumes that an isotropic flux of astrophysical neutrinos is incident on the Earth with a total differential all-flavor neutrino-plus-anti-neutrino spectrum given by
@@
Are the events observed by IceCube consistent with the assumption of SPL? To a large extent, yes. However, there are some tantalizing discrepancies that have motivated a large body of work in the last decade. Some of the explanations are based on dark matter Feldstein et al. (2013); Esmaili and Serpico (2013); Bai et al. (2016); Fong et al. (2015); Rott et al. (2015); Esmaili et al. (2014); Murase et al. (2015); Anchordoqui et al. (2015); Bhattacharya et al. (2017a, b); Hiroshima et al. (2018); Chakravarty et al. (2020), resonances beyond standard model Anchordoqui et al. (2006); Barger and Keung (2013); Dutta et al. (2015); Dey and Mohanty (2016); Dev et al. (2016); Mileo et al. (2016); Chauhan et al. (2018); Dey et al. (2018); Bečirević et al. (2018); Babu et al. (2020); Dey et al. (2021), or non-standard/secret neutrino interactions Ng and Beacom (2014); Ioka and Murase (2014); Ibe and Kaneta (2014); Blum et al. (2014); Araki et al. (2015); Cherry et al. (2014); Kamada and Yu (2015); DiFranzo and Hooper (2015); Araki et al. (2016); Rasmussen et al. (2017); Jeong et al. (2018); Mohanty et al. (2019); Pandey et al. (2019a, b); Bustamante et al. (2020); Creque-Sarbinowski et al. (2021); Mazumdar et al. (2022); Fiorillo et al. (2020a, b); Carpio et al. (2023); Esteban et al. (2021). One such discrepancy is the inferred value of the spectral index in two independent datasets Abbasi et al. (2021a, 2022b); Stettner (2020); Abbasi et al. (2025),
$\gamma_{\texttt{\,astro}}^{\text{\,HESE}}=2.87^{+0.20}_{-0.19}\quad\text{and}\quad\gamma_{\texttt{\,astro}}^{\text{\,Nor. Tr.}}=2.37^{+0.09}_{-0.09}\,.$
The HESE sample selects events (cascades, tracks, and double-cascades) that start inside the instrumented detector volume. Any event that deposits its first significant light inside the fiducial volume, without coinciding hits in the outer region, is considered a starting event. Such a strong veto on the outer layers ensures that the event is likely to be induced by a neutrino rather than a through-going muon. Furthermore, an energy cut-off of 60 TeV is considered to reduce the atmospheric neutrinos from the sample. The Northern Tracks sample selects upward-going muon-like tracks that are mostly through-going muons that originate in neutrino interactions outside the instrumented detector volume. This sample has low energy-threshold and very good direction reconstruction, which also allows for point-source searches with this dataset. The Northern Track events above 100 TeV mostly contain events of astrophysical origin Abbasi et al. (2022b, 2025). If the two samples originate from the same diffuse flux of astrophysical neutrinos, then the $\sim 2.3\sigma$ discrepancy between the two fits to the spectral index is quite puzzling. Interestingly, the Enhanced Starting-Track Event Selection (ESTES) sample, which selects muon-like tracks that start inside the detector irrespective of their direction, reports an intermediate value for the spectral index Abbasi et al. (2024),
$\gamma^{\text{ESTES}}_{\texttt{astro}}=2.58^{+0.10}_{-0.09}.$
We refer the reader to Fig. 7 in Ref. Abbasi et al. (2025) for a compendium of fits to the parameters of astrophysical neutrino flux from various IceCube datasets.
The primary aim of this paper is to answer– How are these fits affected if there are dips in the astrphysical neutrino spectrum? Such dips are a consequence of interaction between astrophysical neutrinos and a cosmic background. If the interaction is through an $s$-channel process, then the absorption feature is narrowly located around the resonance energy $E_{\nu}^{\rm res}=M_{X}^{2}/2m_{b}$, where $M_{X}$ is the mass of the mediator and $m_{b}$ is the mass of the background particle. If the interaction is $t$-channel, then a broad spectrum feature is expected Mohanty et al. (2019). In Ref. Esteban et al. (2021), the effect of non-resonant contribution to the process was found to be significant for larger couplings, and we have included them in our analysis. With only standard model interactions, the dip will occur for neutrinos with energy EeV Weiler (1982); Brdar et al. (2022). However, if there are non-standard/secret interactions mediated by a MeV-scale boson, then the dips can occur in the TeV–PeV range Ng and Beacom (2014). In this paper, we study the scenario proposed in Ref. Chauhan and Mohanty (2018) where the dip is caused by interactions between astrophysical neutrinos and a cosmic sterile-neutrino background.
Figure 1: The test statistic (see Equation 7) for the single power-law fit of 7.5 years HESE sample and Glashow resonance Abbasi et al. (2021a) is shown in blue. In orange, we show the test statistic after including absorption features. The horizontal grey dashed lines are used to denote the 68% C.L. and 90% C.L. intervals. We have shown the results from the 11-year Northern Tracks sample Abbasi et al. (2022b) in green.
@@
$\gamma_{\texttt{\,astro}}^{\text{\,HESE,\,This Work}}=2.60^{+0.20}_{-0.19}$
which is in better agreement with the Northern Tracks sample. As discussed below, our errors are underestimated due to the limitations of our analysis. Using a simplified approach, we also show that the inclusion of these dips does not change the fits to the Northern Tracks significantly. However, for a quantitative assessment of the relaxation in the tension between the two datasets, a detailed Monte Carlo simulation, along with consistent treatment of backgrounds is required, which is beyond the scope of this work. We have shown in detail that an ‘SPL-with-dips’ is a viable fit to the HESE sample and results in a harder spectral index than an SPL-only fit. Our conclusions are also applicable to other origins of the dip(s), for example, non-standard interactions of active neutrinos.
The paper is organized as follows. In Section II, we discuss the methodology followed in the paper. We discuss the model and constraints, the propagation of astrophysical neutrinos, our estimate of the event rates the three datasets (HESE, Northern Tracks, and ESTES), and our likelihood analysis. In Section III, we provide the main results and end with a short summary in Section IV.
@@ II.1 Lagrangian and constraints
In this work, we follow the model of self-interacting sterile neutrino outlined in Ref. Dasgupta and Kopp (2014). The sterile neutrinos are assumed to have self-interactions mediated by a new MeV-scale vector boson. The interaction Lagrangian is given by,
$\mathcal{L}_{\rm int}=g_{X}\bar{\nu}_{s}\gamma_{\mu}\frac{1}{2}(1-\gamma_{5})\nu_{s}X^{\mu}\,$, (2)
where $X^{\mu}$ is the mediator and $g_{X}$ is the self-interaction coupling constant. In this work, we fix the coupling strength as $g_{X}=0.1$, which is sufficiently large for the re-coupling of sterile neutrinos Chu et al. (2015). At energy scales lower than the mediator mass ($M_{X}$), the interaction can be described by a Fermi-like contact interaction with an effective coupling given by $G_{X}=\frac{\sqrt{2}}{8}\frac{g_{X}^{2}}{M_{X}^{2}}$. Cosmological observations like big bang nucleosynthesis (BBN) and cosmic micorwave backgound (CMB) impose stringent constraints on this effective coupling Song et al. (2018). It was shown in Ref. Song et al. (2018) that small values of the effective coupling do not sufficiently delay the production and are thus excluded. The minimum value of $G_{X}$ required to satisfy the BBN constraints at $95\%$ C.L. is $G_{X}\geq 10^{4}G_{F}$. For the mediator mass range relevant to this work, i.e., $M_{X}\sim 10-50$ MeV, $g_{X}>0.01$ is required to satisfy the BBN constraints. We have shown resulting parameter space for both fixed-$g_{X}$ and fixed-$G_{X}$.
@@
The mass-eigenstates and the flavor-eigenstates are related by a $4\times 4$ rotation matrix characterized by six mixing angles. The three mixing angles from the active sector are fixed to their measured values from Ref. Esteban et al. (2020), and the remaining three angles are free parameters of the model. Due to this mixing, all four mass-eigenstates interact via the new mediator, and interactions between propagating mass-eigenstates and background neutrinos are possible Chauhan and Mohanty (2018). In this simplified model, we have five New Physics parameters – three mixing angles ($\theta_{14}\,,\theta_{24}\,,\theta_{34}$), mass of the sterile neutrino ($m_{s}$), and mass of the mediator ($M_{X}$). In a UV-complete model, additional degrees of freedom may be required for generating the mixing and mass of the gauge boson, and for anomaly cancellation.
Constraints from cosmology– The self-interactions of the sterile neutrino delay their production and hence the constraints from BBN on $N_{\rm eff}$ can be avoided. However, the stringest constraints on the model come from the cosmological inference of the sum of neutrino masses Chu et al. (2018). One of the possible ways to avoid this constraint is to assume additional light sterile neutrinos charged under the new gauge symmetry Chu et al. (2018). In this scenario, when these sterile neutrinos recouple, the energy is redistributed among all new species which weakens the constraints on the sum of neutrino mass. In such a scenario, the number density of the cosmological eV-scale sterile neutrino background will be reduced by a factor proportional to the number of additional species. This can be compensated by a small increase in the coupling (as cross-section scales as $g_{X}^{4}$). This possibility was considered in Ref. Chauhan and Mohanty (2018). The other possibility is to consider a fast decay of the sterile neutrino to active neutrinos Chu et al. (2018). As these sterile neutrinos decay after the decoupling, this would result in a new non-thermal distribution of active neutrinos ($f(p)\sim\delta(p-m_{s}/(2\,\text{or}\,3))$) and this may relax the neutrino mass bound Oldengott et al. (2019). The scattering of the cosmogeic neutrinos with the new component would result in similar “dips”, however this intersting possibility requires a separate and dedicated analysis.
Laboratory constraints– In Ref. Fiorillo et al. (2020a), the constrains on self-interacting sterile neutrino from meson decays was evaluated assuming pseudo-scalar interactions. We expect that similar bounds will exist on vector mediators, however, we do not evlauate them. These bounds arise from measurements of leptonic decays of charged mesons ($M\rightarrow\ell+\nu_{\ell}$) which get additional contribution from self-interactions of the form $M\rightarrow\ell+\nu_{\ell}+X$ and $M\rightarrow\ell+\nu_{\ell}+\bar{\nu}_{s}+\nu_{s}$. The four-body decays are supressed, and dominant constraints arise from three-body decays only Fiorillo et al. (2020a). For the light sterile neutrino that is considered in this paper, the bounds on the coupling constant are strong when the new decay modes become kinematically accesible. Based on the results of Ref. Fiorillo et al. (2020a), we expect that our model is strongly constrained for $M_{X}<$ 2 GeV, which is the entire parameter space in consideration. These bounds will be sensitive to details of UV completion, e.g., presence of additional light sterile neutrinos and other mediators. In this respect, the alternative model with a pseudoscalar mediator allows for selective flavor-dependence of the couplings (unlike gauge interactions that are universal) and the limits can be relaxed Blinov et al. (2019). Moreover, the pseudoscalar mediator is also cosmologically more viable tooDasgupta and Kopp (2021).
II.2 Propagation of astrophysical neutrinos
@@
$+\sum_{j,\,k,\,l}n^{t}_{j}\int_{E_{\nu}}^{\infty}dE_{\nu}^{\prime}\,\tilde{n}_{k}(t,E_{\nu}^{\prime})\frac{d\sigma_{jk\rightarrow il}}{dE_{\nu}}(E_{\nu}^{\prime},E_{\nu})\,.$ (3)
@@
In this equation, the first term on the right-hand side accounts for the cosmological expansion. The second term, called the source term, accounts for the production of high-energy neutrinos due to astrophysical sources. In this work, we assume that the TeV–PeV neutrinos originate in $pp$-sources and following Ref. Esteban et al. (2021), the source term is taken to be a single power law with normalization proportional to the star-formation rate. The third term accounts for the interaction with background neutrinos which depends on the assumed number density of sterile neutrinos today. The sterile neutrinos are produced after recoupling from the population of active neutrinos. One can do detailed analysis by solving the 4-flavor Boltzmann transport equation (similar to Refs. Song et al. (2018); Mazumdar et al. (2020)) to obtain the exact number density of the sterile neutrinos. However, we will assume flavor equilibrium for simplicity and assume that the number density of sterile neutrinos ($n_{s}$) can be approximated as $n_{s}=(1/4)\times n_{a}$, where $n_{a}=2\times 56\,(1+z)^{3}$ is the number density of active neutrinos per flavor in the standard scenario. The last term accounts for the regeneration of lower-energy neutrinos after interactions. These equations have been solved using the publicly available code nuSIprop Esteban et al. (2021), which solves the Boltzmann equations for standard three-flavor neutrinos in the presence of non-standard interaction. We have extended the code to include the sterile neutrino. These are improvements on the work done in Ref. Chauhan and Mohanty (2018) where the source term was ad-hoc, and regeneration was ignored.
II.3 High Energy Starting Events (HESE)